# CRAFT: Extracting and Tuning Cultural Instructions from the Wild

### v1

@@ Appendix A Keywords and Templates
In this section, we detail the keywords and instructions employed for selective data extraction, automated question generation, and answer formulation.
@@
For question generation, we employ a straightforward prompt: "You are a chatbot who always generates just one question about Singapore from the given context. Do not generate the answer.". For generating both context-dependent and context-free answers, we use the direct prompt: "Please answer the following question.". The difference lies in whether the context is provided and the model will decide on how much the answer should rely on the provided context.

### v2

@@ 5 Conclusion
Acknowledgement
This work is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG2-GC-2022-005). This research is also supported by the National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore under its National Large Language Models Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority, Singapore.
@@ Appendix A Keywords and Templates
This section details the keywords and instructions employed for selective data extraction, automated question generation, and answer formulation.
@@
For keywords, a complete list can be found in our source code.
@@
We employ a straightforward prompt for question generation: "You are a chatbot who always generates just one question about Singapore from the given context. Do not generate the answer.". For generating both context-dependent and context-free answers, we use the direct prompt: "Please answer the following question.". The difference lies in whether the context is provided and the model will decide on how much the answer should rely on the provided context.